# Adaptive Active Inference Agents for Heterogeneous and Lifelong Federated Learning thanks: *Corresponding author

### v1

@@ Abstract
Handling heterogeneity and unpredictability are two core problems in pervasive computing. The challenge is to seamlessly integrate devices with varying computational resources in a dynamic environment to form a cohesive system that can fulfill the needs of all participants. Existing work on systems that adapt to changing requirements typically focuses on optimizing individual variables or low-level Service Level Objectives (SLOs), such as constraining the usage of specific resources. While low-level control mechanisms permit fine-grained control over a system, they introduce considerable complexity, particularly in dynamic environments. To this end, we propose drawing from Active Inference (AIF), a neuroscientific framework for designing adaptive agents. Specifically, we introduce a conceptual agent for heterogeneous pervasive systems that permits setting global systems constraints as high-level SLOs. Instead of manually setting low-level SLOs, the system finds an equilibrium that can adapt to environmental changes. We demonstrate the viability of AIF agents with an extensive experiment design, using heterogeneous and lifelong federated learning as an application scenario. We conduct our experiments on a physical testbed of devices with different resource types and vendor specifications. The results provide convincing evidence that an AIF agent can adapt a system to environmental changes. In particular, the AIF agent can balance competing SLOs in resource heterogeneous environments to ensure up to 98% fulfillment rate.
@@
The Distributed Computing Continuum is an emerging paradigm for systems that can seamlessly integrate multiple layers of computing infrastructure [1]. Computing continuum systems promise to enable infrastructure-critical pervasive applications with stringent requirements, such as Mobile Augmented Reality (MAR) for cognitive applications [2] and Remote Sensing for Disaster Management [3]. There are three recurrent characteristics among pervasive applications deployed on a continuum. First is their reliance on AI-based methods for tasks that classical control structures cannot solve efficiently or with sufficient precision [4]. For example, MAR applications must process streams of high-dimensional data that a service could ideally process at the source to fulfill a sub-10 millisecond latency Service Level Objective (SLO). The caveat is that resources in proximity are constrained. Typical solutions involve task partitioning and lightweight data reduction methods that minimize the penalty for offloading to remote resources [5]. The second is heterogeneity, i.e., resource-asymmetry, vendor specifications, and usage patterns. Although pervasive applications follow an overall common objective, a system must consider the individual properties and objectives of participants. The third is the continuously drifting problem domain intrinsic to the dynamic environments of pervasive applications, such that the source distribution drifts over time and data volume is non-static. Conclusively, a necessary precondition is a system that can adapt to non-identically and independently distributed (non-IID) data. Moreover, the system must facilitate collaboration between heterogeneous devices to fulfill their SLOs by distributing workload fairly and considering the individual properties of participants.
@@
The focus is on lifelong heterogeneous federated learning (FL) as we find it best encapsulates the primary challenges of pervasive applications that share the described characteristics. In general, FL participants collaborate for a common objective, i.e., to maximize the prediction performance. Yet, each participant has a private local validation set to determine whether their criteria are locally met. Time constraints that ensure smooth operations and resource asymmetry further instigate friction when attempting to satisfy local objectives. Hence, despite a common objective, to fulfill the SLOs of each participant individually, a delicate balance is necessary. Lastly, the dynamic environment gradually drifts the distribution and varies the data volume.
@@
This work aims to demonstrate the viability of Active Inference (AIF) in designing adaptive agents that can gracefully handle the challenging requirements of pervasive applications. While Active Inference is a neuroscientific framework, recent work has shown promising results by conceiving methods from the underlying ideas for workload scheduling in distributed systems. In particular, we find that the objectives of Active Inference and pervasive computing intrinsically intertwine. Context awareness is crucial for pervasive applications as these systems operate in dynamic environments and must adapt to changes in their surroundings. Precisely context awareness is a defining characteristic of AIF agents. However, the current application of AIF for systems is more conceptual and only partially implements the core components of the AIF framework [6, 7]. Other work on systems that adapt to changing requirements typically focuses on optimizing individual variables, such as learning rate or setting low-level SLOs as constraints on specific resources [8, 9]. Despite providing more fine-grained control, it is unreasonable to expect application developers to understand the implications of each low-level constraint to the overall system, particularly in dynamic environments where resources are scarce and availability is less predictable. In contrast, our AIF agent permits setting high-level SLO targets to find an equilibrium without attempting to enforce constraints from possibly conflicting low-level SLOs.
@@
  • A conceptual AIF agent that balances multiple SLOs during model training. When SLOs compete, agents can autonomously infer optimal training configurations without manual intervention.
  • The empirical evaluation of AIF agents for pervasive FL tasks; experiments were designed to reflect real-world conditions, such as data and resource heterogeneity
II Background
@@ II-A Lifelong Heterogeneous Federated Learning
II-A1 Heterogeneous Federated Learning
We refer to Federated Learning as heterogeneous when data is non-IID and hardware specifications vary among participants. Moreover, hardware heterogeneity typically implies that resources are constrained, i.e., to not discriminate against computationally less powerful devices closer to the source.
II-A2 Lifelong Federated Learning
FL is lifelong when training continuously adapts to concept drifts  [10]. Introducing concept drifts is an intrinsic property of the dynamic deployment environment of pervasive applications. Yet, there is limited research interest in lifelong learning for FL [11].
@@ II-A3 Service Level Objectives for Federated Learning
For our purposes, high-level SLOs provide an intuitive interface to set targets for an AIF agent and quantitative measures to determine its adaptability to a dynamic environment. In particular, maintaining prediction performance and minimizing round duration are two primary objectives for lifelong heterogeneous federated learning. An SLO on prediction performance ensures consistent solution quality. In contrast, an SLO on timeliness is crucial as resources are constrained, and a considerably slower client can delay global weight updates.
@@ II-B Active Inference
AIF agents continuously evaluate the expected free energy (EFE) for different policies and assess their impact on underlying models. In a system’s context, an agent who understands the world model will minimize EFE by selecting policies likely to fulfill SLOs.
We can describe the EFE with two distinct components [16]:
@@
The information gain (IG) estimates how much the model can improve by choosing a particular policy. Conversely, the pragmatic value estimates how close a possible outcome will be to the agent’s preferred state.
@@ III Problem Statement
We consider a lifelong heterogeneous FL system consisting of an orchestrator with $N$ participants. Figure 1 illustrates an example.
@@
Client hardware is heterogeneous in vendor specification, available resource types, and overall computational capacity. For example, some devices may have onboard accelerators, such as GPUs, and others may only work with energy-efficient CPUs. The data source is non-IID with temporal correlations, i.e., the training must adapt to non-stationary data. SLOs are globally configurable, and clients check after each training round whether the SLOs are fulfilled locally. The SLOs aim to ensure smooth operations, i.e., timely training and consistently adequate model performance. We introduce a mechanism to control and manage SLO fulfillment by defining FL training configurations, which specify training parameters that directly affect the system’s ability to fulfill SLOs. Configurations function as levers that an agent can change to fulfill the user-set high-level SLOs. The objective is for the FL system to maximize the overall SLO fulfillment across all timestamps. The challenge is to ensure the highest possible SLO fulfillment, given heterogeneous participants and data within a dynamic environment.
The following describes how we can draw from AIF to design agents that can result in a self-adapting SLO-aware system behavior. In short, the agent learns the world model and becomes sensitive to changes, such that it can maintain service quality by reacting to environmental changes.
@@ IV Proposed Method
interact. Algorithm 1 summarizes the overall procedure for client-side training.
@@
The following elaborates on notable details about the process for an agent to find and choose optimal FL training configurations in a dynamic environment.
IV-A Learning a Simple Causal World Model
The generative model is at the core of an AIF agent, i.e., as an agent interacts with its environment, it updates its internal world representation in a perception-action cycle. We choose Bayesian Networks (BNs) as they provide interpretable graphical representations of learned causal structures. This work considers discrete BNs with uniform priors. Each agent’s initial Bayesian network structure is unknown, as there are no assumptions on prior knowledge of the environmental dynamic. The agents require only starting knowledge of the BN variables and their respective cardinalities. We define the BN $\mathcal{B}$ of an agent as:
@@ IV-A Learning a Simple Causal World Model
We use Hill Climb Search [17] and Bayesian estimation to perform structural learning and parameter estimation. We use variable estimation to perform exact inference, such that an AIF agent utilizes precise computation to leverage the uncertainty of BNs. As the FL rounds progress, the BN causal structures are progressively discovered. Moreover, the agent experiences further refine the conditional probability distributions.
@@ IV-B EFE and SLO-aware Configuration Selection
Due to SLO and configuration vertices present in the BN, the task is to correctly discover the connections between the vertices to choose the configuration with the highest chance of fulfilling SLOs. The agent calculates the EFE for each configuration available to the system to determine configuration optimality using the formula in Equation 1. Specifically, it calculates the pragmatic value as:
@@
The preference vector encodes the agent’s goal, i.e., the desired outcome, and is expressed as a logarithm of the normalized preferences. The list of outcomes consists of all combinations of possible SLOs values. Since the agent is to fulfill all set SLOs, the state where all SLOs are fulfilled will have a higher preference. For example, for one binary outcome SLO, one can set the preference vector to [0.001, 0.999], specifying that the second outcome is preferred. Information gain quantifies the expected Bayesian surprise that measures how much observing new data would update the agent’s belief about hidden states. Following the implementation in [18], the calculation uses the likelihood of SLOs fulfillment under configurations (matrix $A$) and the predictive density over hidden states $q$ derived from the Bayesian Network. An observation is a configuration and an associated outcome. The agent uses a particular configuration as evidence to predict the possible observation with a maximum a posteriori (MAP) query to simulate possible future and predict its IG. In summary, the information gain and pragmatic value balance a trade-off between exploring and taking the actions that most likely result in SLOs fulfillment. Once the agent has selected a configuration, the local training round starts. On completion, the agent collects lower-level metrics and checks SLO fulfillment. Lastly, it associates the outcomes with the configuration and adds it to the history dataset as a new observation for further updates. A BN update step can result in a parameter update or structure learning followed by a parameter estimation. Figure 3 illustrates the process of learning the structure of the BN as the FL rounds progress.
We distinguish between two update types to allow the agent to adapt to significant discrepancies between expected and observed outcomes. If the observed IG is higher than expected, the agent discards structure information from the previous iterations and initializes structure re-learning of the BN. Structure re-learning prioritizes the edges that include SLOs as the dependent node, to ensure that the relations that have an immediate impact on agent decisions are considered first. Conversely, if the observed IG is within expectations, the agent only initializes a parameter update on the BN.
To ensure the BN does not learn from early-stage ML model performance data, that do not accurately describe the relationship between performance SLO and configuration, these observations are omitted until the model performance stabilizes. When the global FL model performance gets sufficiently close to the target performance SLO, a lifelong learning flag signals the AIF agent to start learning.

### v2

@@ Abstract
Handling heterogeneity and unpredictability are two core problems in pervasive computing. The challenge is to seamlessly integrate devices with varying computational resources in a dynamic environment to form a cohesive system that can fulfill the needs of all participants. Existing work on adaptive systems typically focuses on optimizing individual variables or low-level Service Level Objectives (SLOs), such as constraining the usage of specific resources. While low-level control mechanisms permit fine-grained control over a system, they introduce considerable complexity, particularly in dynamic environments. To this end, we propose drawing from Active Inference (AIF), a neuroscientific framework for designing adaptive agents. Specifically, we introduce a conceptual agent for heterogeneous pervasive systems that permits setting global systems constraints as high-level SLOs. Instead of manually setting low-level SLOs, the system finds an equilibrium that can adapt to environmental changes. We demonstrate the viability of our AIF agents with an extensive experiment design, using heterogeneous and lifelong federated learning as an application scenario. We conduct our experiments on a physical testbed of devices with different resource types and vendor specifications. The results provide convincing evidence that an AIF agent can adapt a system to environmental changes. In particular, the AIF agent can balance competing SLOs in resource heterogeneous environments to ensure up to 98% fulfillment rate.
@@
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@
The Distributed Computing Continuum is an emerging paradigm for systems that can seamlessly integrate multiple layers of computing infrastructure [1]. Computing continuum systems promise to enable infrastructure-critical pervasive applications with stringent requirements, such as Mobile Augmented Reality (MAR) for cognitive applications [2] and remote sensing for disaster management [3]. There are three recurrent characteristics among pervasive applications deployed on a continuum. First, is their reliance on AI-based methods for tasks that classical control structures cannot solve efficiently or with sufficient precision [4]. For example, MAR applications must process streams of high-dimensional data that a service could ideally process at the source to fulfill a sub-10 millisecond latency Service Level Objective (SLO). The caveat is that resources in proximity are constrained. Typical solutions involve task partitioning and lightweight data reduction methods that minimize the penalty for offloading to remote resources [5]. The second characteristic is heterogeneity, i.e., resource-asymmetry, vendor specifications, and usage patterns. Although pervasive applications follow an overall common objective, a system must consider the individual properties and objectives of participants. Third, is the continuously drifting problem domain intrinsic to the dynamic environments of pervasive applications, such that the source distribution drifts over time and data volume is non-static. Conclusively, a necessary precondition is a system that can adapt to non-identically and independently distributed (non-IID) data. Moreover, the system must facilitate collaboration between heterogeneous devices to fulfill their SLOs by distributing workload fairly and considering the individual properties of participants.
@@
The focus of this paper is on lifelong heterogeneous federated learning (FL) as we find it best encapsulates the primary challenges of pervasive applications that share the described characteristics. In general, FL participants collaborate for a common objective, i.e., to maximize the prediction performance. Yet, each participant has a private local validation set to determine whether their criteria are locally met. Time constraints that ensure smooth operations and resource asymmetry further instigate friction when attempting to satisfy local objectives. Hence, despite a common objective, to fulfill the SLOs of each participant individually, a delicate balance is necessary. Lastly, the dynamic environment gradually drifts the distribution and varies the data volume.
@@
This work aims to demonstrate the viability of Active Inference (AIF) in designing adaptive agents that can gracefully handle the challenging requirements of pervasive applications. While AIF is a neuroscientific framework, recent work has shown promising results by conceiving methods from the underlying ideas for workload scheduling in distributed systems [6, 7]. In particular, we find that the objectives of Active Inference and pervasive computing intrinsically intertwine. Context awareness is crucial for pervasive applications as these systems operate in dynamic environments and must adapt to changes in their surroundings. Precisely context awareness is a defining characteristic of AIF agents. However, the current application of AIF for systems is more conceptual and only partially implements the core components of the AIF framework. Other work on systems that adapt to changing requirements typically focuses on optimizing individual variables, such as learning rate or setting low-level SLOs as constraints on specific resources [8, 9]. Despite providing more fine-grained control, it is unreasonable to expect application developers to understand the implications of each low-level constraint to the overall system, particularly in dynamic environments where resources are scarce and availability is less predictable. In contrast, our AIF agent permits setting high-level SLO targets to find an equilibrium, defined as the system configuration that fulfills all its SLOs, without attempting to enforce constraints from possibly conflicting low-level SLOs.
@@
  • A conceptual AIF agent that balances multiple SLOs during model training. When SLOs have competing targets, agents can autonomously infer optimal training configurations without manual intervention.
  • Empirical evaluation of AIF agents for pervasive FL tasks under real-world-inspired conditions, incorporating data and resource heterogeneity through a reproducible experimental setup with a physical testbed of heterogeneous devices.
II Background and Related Work
@@ II-A Lifelong Heterogeneous Federated Learning
II-A1 Lifelong Federated Learning
FL is lifelong when training continuously adapts to concept drifts and other changes occurring in continuous data streams  [10]. Introducing concept drifts is an intrinsic property of the dynamic deployment environment of pervasive applications. The presence of concept drift can lead to both prolonged time until model convergence and reduced model performance, stressing the importance of treating concept drifts in FL and the need for targeted solutions for different types of concept drifts [11]. Yet, there is limited research in lifelong learning for FL [12]. The current approaches to adapt to concept drifts rely on custom drift detectors to understand when the drift occurs [13, 14], which results in the need to tune said detectors to the use case and potential drift scenarios, leaving the challenge on how to distinguish drifts from anomalous data. Moreover, although concept drifts are one of the most challenging issues to face for lifelong FL, they are not the only one, as discussed further in this section. This highlights the need for more self-adapting mechanisms, e.g., by introducing meta-learning concepts to FL [15], that can maintain the performance of the FL systems.
II-A2 Heterogeneous Federated Learning
We refer to Federated Learning as heterogeneous when data is non-IID and hardware specifications vary among participants. Moreover, hardware heterogeneity typically implies that resources are constrained, as less powerful devices, often located closer to the data source, must also be accommodated in the learning process.
Optimizing FL workflows, especially in the presence of heterogeneities, is important for minimizing the time-to-accuracy of model training [16].
In that context, Kundroo et al. [9] proposed FedHPO, a federated optimization algorithm that accelerates each client’s training by modifying its hyperparameters, such as learning rate or epochs. However, FedHPO introduces additional algorithm parameters to set and tune, e.g., patience or thresholds to guide the optimization process, which limits its flexibility in dynamic environment usage. To optimize for local training time in dynamic and heterogeneous devices conditions, an asynchronous FL approach FedTS was proposed by Li et al. [17], empowering the FL server to detect and optimize for slower clients. Still, the scheme focuses on ensuring time constraints for heterogeneous devices and does not cover lifelong scenarios with dynamic data distributions.
Several studies also explored multi-objective optimization (MOO) in FL to balance competing objectives. One approach is to optimize neural network models instead of client-specific hyperparameter optimization [18, 19]. Additionally, a significant number of existing research focuses on optimizing client selection or clustering instead of adjusting the parameters of individual clients to reach multi-dimensional goals [20, 21, 22]. Although these approaches offer practical solutions for balanced MOO, they do not take into account the individual parameters of heterogeneous clients. Moreover, a vast area of research lies in adopting Bayesian Optimization (BO) [23] to the needs of FL. Along with grid search, BO is used for hyperparameter tuning in FL [24], still, the downside of both methods includes their inherent incentive to find the best possible hyperparameters set (either in one-shot fashion or during multiple communication rounds) which makes it more difficult to adapt them to changing conditions of lifelong learning. Another drawback of the classical BO is its focus on one singular objective to optimize, e.g., model accuracy. This problem is starting to be addressed by Multi-objective Bayesian Optimization (MOBO) [20]. Yet, the presented MOBO approach also does not consider lifelong learning scenarios.
A number of automated optimization tools are available to use for hyperparameter optimization tasks and were adopted by the industry, such as HyperOpt [25] and Optuna [26], with such platforms as RayTune [27] allowing for distributed parameter tuning. Lately, FL research started to adopt hyperparameter optimization tools [28, 29, 30, 31, 32]. However, currently, the application of hyperparameter tuning still does not cover lifelong learning scenarios.
@@ II-A3 Service Level Objectives for Federated Learning
For our purposes, high-level SLOs provide an intuitive interface to set targets for an AIF agent and quantitative measures to determine its adaptability to a dynamic environment. In particular, maintaining prediction performance and minimizing round duration are two primary objectives for lifelong heterogeneous federated learning. An SLO on prediction performance ensures consistent solution quality. In contrast, an SLO on timeliness is crucial as resources are constrained, and a considerably slower client can delay global weight updates. Time and prediction performance SLOs abstract more detailed system parameters that have an impact on them, focusing on end-user experience and overall system performance.
There exist multiple approaches that aim to combine SLOs with dynamic processing requirements: Zhang et al. [35] presented Octopus – the framework that finds optimal services configurations in multi-tenant edge computing scenarios. Octopus predicts SLO fulfillment of two variables based on a deep neural network. Shubha et al. [8] presented AdaInf, which detects SLO violations of a GPU scheduling task whenever variable drifts occur. Through AdaInf, it is possible to find SLO-fulfilling resource allocations between model training and inference. Although these approaches are SLO-aware, oriented at continuous processes, and may utilize agents for the decision-making process, they are primarily used for inference services and do not consider the scenario of optimizing the training of ML models, which require more flexible and high-level SLOs definitions, appropriate for the considered FL/ML scenario.
@@ II-B Active Inference
AIF agents continuously evaluate the expected free energy (EFE) for different policies and assess their impact on underlying models. EFE constitutes the planning ability of AIF agents, as it allows for evaluating policies of custom length into the future, utilizing the current understanding of the world model (generative world model) as the source for simulations. In a system’s context, an agent who understands the world model will minimize EFE by selecting policies likely to fulfill SLOs. We describe EFE with two distinct components [38]:
@@
The information gain (IG) estimates how much the model can improve by choosing a particular policy and aims to resolve the uncertainty currently present in the generative world model. Thus, IG takes into consideration the approximate posterior Q over hidden states (variables in the world that the agent cannot observe) given observations and Q of hidden states only – our prior beliefs about the model before any observation. Here, the agent aims at maximizing the divergence, thus, looking for the most informative future steps. Conversely, the pragmatic value (PV) estimates how close a possible outcome (observation O) will be to the agent’s preferred observations (C), focusing on meeting the expectation defined for the agent. Together, IG and PV balance the exploration/exploitation trade-off.
To sum up, while AdaInf and Octopus are SLO-aware inference services, they are designed for dynamic and resource-efficient model serving, particularly in scenarios like multi-model and edge inference. In contrast, traditional FL hyperparameter tuning methods, such as BO and grid search, aim to achieve one best FL model but do not inherently address the adaptability and continuity required for lifelong learning. Active Inference, with its inherent adaptability, innate exploration-exploitation trade-off handling, and potential explainability through causality, offers a promising theoretical solution for the presented gap. To the best of our knowledge, this paper is the first attempt to apply AIF to FL (and, by extension, lifelong FL), addressing the challenges of adaptability and heterogeneity in this domain.
@@ III Problem Statement
We consider a lifelong heterogeneous FL system consisting of an orchestrator with $N$ participants. In Figure 1 we can observe a central model and orchestrator in the Cloud that communicates with different IoT/Edge devices. These produce a stream of data that is used by an ML model; at the same time, these devices embed an AIF agent that adjusts the training of their ML models to achieve optimal (SLO fulfillment) performance.
@@
The objective of the FL system described in Figure 1 is to maximize the overall SLO fulfillment across all timestamps. The challenge is to ensure the highest possible SLO fulfillment (reaching equilibrium), given heterogeneous participants and data within a dynamic environment. Client hardware is heterogeneous in vendor specification, available resource types, and overall computational capacity. For example, some devices may have onboard accelerators, such as GPUs, and others may only work with energy-efficient CPUs. The data source is non-IID with temporal correlations, i.e., the training must adapt to non-stationary data. SLOs are set on a global system level and all clients aim to fulfill the same SLOs. SLOs aim to ensure smooth operations, i.e., timely training and consistently adequate model performance, and clients check after each training round whether the SLOs are fulfilled locally.
We introduce a mechanism to control and manage SLO fulfillment by defining FL training configurations, which specify training parameters that directly affect the system’s ability to fulfill SLOs. Configurations function as levers that an agent can change to fulfill the high-level SLOs.
@@ IV Proposed Method
interact. (1) The orchestrator sends the current global model to the client, initiating a new FL round; (2) the client requests a configuration from the AIF Agent; (3) the AIF agent evaluates configurations based on EFE and returns the best one; (4) the client trains the model using the selected configuration; (5) after training, the client sends a new observation to the AIF Agent and returns the updated model to the Orchestrator, which aggregates all updates; (6) the AIF agent updates its world model (Bayesian Network) for future configuration optimization; (7) the next FL round begins. Algorithm 1 summarizes the overall procedure for client-side training with the rest of the section elaborating on notable details about the process for an agent to find and choose optimal FL training configurations in a dynamic environment.
@@
IV-A Learning a Simple World Model
The generative model is at the core of an AIF agent, i.e., as an agent interacts with its environment, it updates its internal world representation in a perception-action cycle to improve its understanding and align its behavior to reach set goals. We choose Bayesian Networks (BNs) as they provide interpretable graphical representations of learned causal structures and offer a principled framework for probabilistic reasoning. This work considers discrete BNs with uniform priors. Each agent’s initial Bayesian network structure is unknown, as there are no assumptions on prior knowledge of the environmental dynamic. The agents require only starting knowledge of the BN variables and their respective cardinalities, thus, specifying the considered features in the environment and their respective precision. We define the BN $\mathcal{B}$ of an agent as:
@@ IV-A Learning a Simple World Model
Figure 3 illustrates the process of learning the structure of the BN as the FL rounds progress. We use Hill Climb Search [40] and Bayesian estimation to perform structural learning and parameter estimation. We use variable estimation to perform exact inference, such that an AIF agent utilizes precise computation to leverage the uncertainty of BNs. As the FL rounds progress, the BN causal structures are progressively discovered. Moreover, the agent experiences further refine the conditional probability distributions.
To allow the agent to adapt to significant discrepancies between expected and observed outcomes, we distinguish between two update types. If the observed IG is higher than expected, the agent discards structure information from the previous iterations and initializes structure re-learning of the BN. Structure re-learning prioritizes the edges that include SLOs as the dependent node, to ensure that the relations that have an immediate impact on agent decisions are considered first. Conversely, if the observed IG is within expectations (equal or below expected IG), the agent only initializes a parameter update on the BN, merely accommodating new data that followed the expectations of the current generative model.
To ensure the BN does not learn from early-stage ML model performance data, that do not accurately describe the relationship between performance SLO and configuration, and instead focuses on the lifelong part of the training only, observations from early FL rounds are omitted until the model performance stabilizes. When the global FL model performance gets sufficiently close to the target performance SLO, a lifelong learning flag signals the AIF agent to start learning, before that, the configuration is taken at random, as no pre-defined strategies are defined for this warm-up stage.
@@ IV-B EFE and SLO-aware Configuration Selection
Due to SLO and configuration vertices present in the BN, the task is to choose the configuration with the highest chance of fulfilling SLOs by correctly discovering the connections between the vertices and resolving uncertainty about the world. The agent calculates the EFE for each configuration available to the system to determine configuration optimality using the formula in Equation 1. Specifically, it calculates the pragmatic value as:
@@
The preference vector encodes the agent’s goal, i.e., the desired outcome, and is expressed as a logarithm of the normalized preferences. The list of outcomes consists of all combinations of possible SLO values. Since the agent is to fulfill all set SLOs, the state where all SLOs are fulfilled will have a higher preference. For example, for one binary outcome SLO, one can set the preference vector to [0.001, 0.999], specifying that the second outcome is preferred.
Information gain quantifies the expected Bayesian surprise that measures how much observing new data would update the agent’s belief about hidden states. First, for each possible configuration, the agent assesses which observation is most likely. An observation is a configuration and an associated outcome. The agent uses a particular configuration as evidence to predict the possible observation with a maximum a posteriori (MAP) query to simulate possible future. Then, following the implementation in [41], the calculation (Equation 3) uses the likelihood of SLOs fulfillment (matrix $A$) and the predictive density over hidden states $q$ derived from the BN to predict IG of a specific configuration.
In summary, the information gain and pragmatic value balance a trade-off between exploring and taking the actions that most likely result in SLOs fulfillment. Once the agent has selected a configuration, the local training round starts. On completion, the agent collects lower-level metrics and checks SLO fulfillment. Lastly, it associates the outcomes with the configuration and adds it to the history dataset as a new observation for further updates.